# FRRI: a novel algorithm for fuzzy-rough rule induction

### v1

@@ 4.2 Results
FRRI MODLEM QuickRules FURIA RIPPER
FRRI 1.00e+0 4.07e-1 9.41e-3 5.13e-2 7.31e-4
MODLEM 4.07e-1 1.00e+0 5.74e-2 6.98e-3 4.03e-5
QuickRules 9.41e-3 5.74e-2 1.00e+0 2.00e-6 3.23e-9
FURIA 5.13e-2 6.98e-3 1.77e-6 1.00e+0 2.44e-1
RIPPER 7.31e-4 4.00e-5 3.23e-9 2.44e-1 1.00e+0
@@
Let us now look at the average size of the rulesets generated by the algorithms, where the Friedman test indicates the presence of significant ($p<0.0001$) differences between the algorithms. Table 6 contains the $p$-values obtained by performing the Conover post-hoc pairwise comparison procedure. First, we notice that FURIA and RIPPER generate significantly smaller rulesets than the (fuzzy)-rough rule induction algorithms. This is because both of these models use an iterative pruning process, where old rules are repeatedly replaced with more general versions. Moreover, QuickRules is by far the worst performer. One reason this might be the case, is that the iterative process of QuickRules considers the objects in the training set many times, and looks for a total ruleset that covers the entire dataset in a greedy way. Without any pruning, this is sure to deliver much more rules that algorithms with better heuristics (MODLEM), pruning (FURIA and RIPPER) and those that use the exact solutions of an optimization problem (FRRI). Finally, let us zoom in on the contrast between FRRI and MODLEM, which is even more pronounced here. FRRI generates the smaller ruleset of the two in all but 5 of our datasets. It is interesting to observe that those datasets are highly imbalanced. Actually, this maybe considered further evidence that FRRI is better suited towards dealing with imbalanced data, as it will generate enough rules to cover both the smallest and biggest classes completely. For MODLEM, the worst case is vowel, with where it needs 77 more rules. On the other hand, statistical testing again shows that both MODLEM and FRRI are significantly better ($p<0.001$) than QuickRules. Using a one-sided Wilcoxon test, we can establish that FRRI is also significantly better ($p<0.05$) than MODLEM w.r.t. the size of the induced ruleset.
@@ 5 Conclusions and future work
In this article, we introduced a novel rule induction algorithm, Fuzzy-Rough Rule Induction (FRRI), which combines the best parts of fuzzy and rough set theory to induce small rulesets consisting of high-accuracy rules. Our algorithm starts by discarding unnecessary information from the objects in our dataset, and then selects a minimal amount of objects to serve as rules. These rules can be viewed as summaries of the data, which are put to work to classify new samples. Experimental evaluation of the performance of FRRI against state-of-the-art rule induction algorithms in this last task showed that our algorithm is more accurate while using fewer rules. Moreover, the obtained rules are not much longer than those induced by state-of-the-art rough set algorithms.

### v2

@@ 4.2 Results
FRRI MODLEM QuickRules FURIA
MODLEM $0.407$
QuickRules $0.009$ $0.057$
FURIA $0.051$ $0.007$ $<0.001$
RIPPER $<0.001$ $<0.001$ $<0.001$ $0.244$
@@
Let us now look at the average size of the rulesets generated by the algorithms, where the Friedman test indicates the presence of significant ($p<0.0001$) differences between the algorithms. Table 6 contains the $p$-values obtained by performing the Conover post-hoc pairwise comparison procedure. First, we notice that FURIA and RIPPER generate significantly smaller rulesets than the (fuzzy)-rough rule induction algorithms. This is because both of these models use an iterative pruning process, where old rules are repeatedly replaced with more general versions. Moreover, QuickRules is by far the worst performer. One reason this might be the case, is that the iterative process of QuickRules considers the objects in the training set many times, and looks for a total ruleset that covers the entire dataset in a greedy way. Without any pruning, this is sure to deliver much more rules that algorithms with better heuristics (MODLEM), pruning (FURIA and RIPPER) and those that use the exact solutions of an optimisation problem (FRRI). Finally, let us zoom in on the contrast between FRRI and MODLEM, which is even more pronounced here. FRRI generates the smaller ruleset of the two in all but 5 of our datasets. It is interesting to observe that those datasets are highly imbalanced. Actually, this maybe considered further evidence that FRRI is better suited towards dealing with imbalanced data, as it will generate enough rules to cover both the smallest and biggest classes completely. For MODLEM, the worst case is vowel, with where it needs 77 more rules. On the other hand, statistical testing again shows that both MODLEM and FRRI are significantly better ($p<0.001$) than QuickRules. Using a one-sided Wilcoxon test, we can establish that FRRI is also significantly better ($p<0.05$) than MODLEM w.r.t. the size of the induced ruleset.
@@ 5 Conclusions and future work
In this article, we introduced a novel rule induction algorithm, Fuzzy-Rough Rule Induction (FRRI), which combines the best parts of fuzzy and rough set theory to induce small rulesets consisting of high-accuracy rules. Our algorithm starts by discarding unnecessary information from the objects in our dataset, and then selects a minimal amount of objects to serve as rules. These rules can be viewed as summaries of the data, which are put to work to classify new samples. Experimental evaluation of the performance of FRRI in this last task showed that our algorithm is on average more accurate than state-of-the-art rule induction algorithms while using fewer rules than those that do not apply iterative pruning approaches. Moreover, the obtained rules are not much longer than those induced by state-of-the-art rough set algorithms.